OpenStudy (anonymous):

evaluate: dx over the square root of 28-12 x^2 ?

OpenStudy (callisto):

Do you mean \[\int \frac{dx}{\sqrt{28-12x^2}}\]for the question?

OpenStudy (anonymous):

yes :)

myininaya (myininaya):

\[\int\limits_{}^{}\frac{dx}{\sqrt{28} \sqrt{1-\frac{12x^2}{28}}} \] \[\frac{1}{\sqrt{28}} \int\limits_{}^{}\frac{dx}{\sqrt{1-(\frac{\sqrt{12} x}{\sqrt{28}}})^2 }\]

myininaya (myininaya):

\[u=\frac{\sqrt{12}x}{\sqrt{28}}\]

OpenStudy (anonymous):

how about its du; dv; and v ?

OpenStudy (anonymous):

integration by parts

OpenStudy (callisto):

I did it using trigo sub...

OpenStudy (anonymous):

integration by parts, (callisto)

OpenStudy (callisto):

Must you use by parts?

OpenStudy (anonymous):

∫112(2812−x2)−−−−−−−−−−√dx=>112−−√∫12812−x2−−−−−−√dx how did you come up with this solution?

OpenStudy (anonymous):

#callisto: yes, it's by parts

OpenStudy (callisto):

Wait, I did something wrong there :|

OpenStudy (anonymous):

its okey, no need to worry

OpenStudy (callisto):

I hope I did it right this time, though, I'm not using integration by parts to get it